# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v2

@@ 4.3. Cross-paradigm Insights and Suggestions
The model’s information-flow structure explains why interaction timing and artifact availability form the principal axes that differentiate the twelve paradigms; the paradigms themselves illustrate how recurring graph topologies map to distinct design affordances. In addition, our model disambiguates interface cases that prior frameworks conflate. The motivating examples in Fig. 2 (b–f) map cleanly to paradigm graphs P1, P3, P4, P7, and P9, respectively, as discussed above, demonstrating meaningful descriptive and discriminative power of the model. Building on this analytical foundation, we derive four concrete, model-grounded considerations that help designers reason about which paradigm class aligns with a given design situation based on interaction timing (C1) and artifact availability (C2), which specific paradigm within that class best matches user intent expression needs, as shown in Tab. 3 (C3), and how the IAI model and derived paradigms can be used generatively to propose new interfaces (C4).
@@
C2. Use artifact availability (prompt-only vs. artifact-grounded) to select how intent should be expressed. Instructions in natural language alone can be ambiguous (e.g., make the flower brighter” without specifying which flower). The presence of an artifact shifts intent expression from purely linguistic descriptions to concrete, grounded references   (Tab. 3 P4, P9-P12). By enabling users to select, highlight, or annotate artifact fragments, systems can reduce ambiguity and provide a stable foundation for commands (Singh et al., 2024; Chen et al., 2023). The consideration also aligns with design principles for AI-instruments by Riche et al.  (Riche et al., 2025), where they proposed that text prompts should be grounded in other artifacts.   Designers should identify whether their system will receive an initial artifact; if so, artifact-grounded paradigms offer lower ambiguity, richer referential precision, and reduced prompting burden.
@@
C3. Within a chosen class, refine toward a specific paradigm based on how intent should be encoded. In these design paradigms, a key design choice is how user intent is materialized to augment the original text prompt. The first way is to directly edit or extend the text prompt with interactions (Tab. 2-P1 and P5). An example is DreamSheets (Almeda et al., 2024), where users can combine prompt fragments easily with interactions. It provides a simple and direct way to improve original prompts but lacks the power to cater complex intent, such as non-linguistic or referential intent. To handle more complex intent, an approach is to leverage interactions to introduce additional non-linguistic information to text and form augmented instructions, such as structural or parametric information (Tab. 2-P2, P3, P6, P7, P9 and P11). For example, CoPrompt (Feng et al., 2024) allows users to drag and drop prompt fragments to form a multi-level list of instructions. It materializes the non-linguistic intent expression in a convenient way with suitable interaction design. Regarding the referential intent, the key interaction is to link text prompts with artifacts and generate augmented instructions (Tab. 2-P4, P10). It addresses the challenges in describing the linkage between texts and artifacts by materializing users’ intent with direct interactions on artifacts. Lastly, a unique case is to generate new interaction widget by text prompts and facilitate intent expression through direct manipulation (Tab. 2-P8, P12). GenAI creates reusable and persistent interactive widgets to materialize the meta intent and allows follow-up similar intent expression with simple interactions with widgets. A notable example is DynaVis (Vaithilingam et al., 2024). The meta intent like changing visual element colors can be materialized as a widget. The specific color to use can be directly selected by interactions.   The summarized design cheat sheet mapping common design situations to recommended paradigms is shown in Tab. 3. These differences correspond to distinct subgraph configurations in the atomic paradigm graphs. Designers can therefore match their intent-expression needs (e.g., decomposition, parameterization, selection specificity) to the paradigm whose graph topology supports that form of control.
@@
Table 4.   Operationalizing Generative Power: Potential methods for designers and researchers to apply the IAI model and derived paradigms. In this table, starting points refer to whether the designer or researcher should look into IAI model or our summarized paradigms at the beginning of the usage method.
@@
C4. Reusing, chaining, and innovating atomic paradigms for adapting interaction design to new scenarios. Our summarized paradigms and the considerations C1-C3 above provide a concise mapping from task requirements to interface affordances.   Designers can start from intent clarity and artifact availability to select appropriate paradigms for interface implementation. For example, they can use the pre-invocation, artifact-grounded paradigm   (Tab. 3 P5) for tasks that start from existing artifacts and demand precision. They can apply post-invocation, prompt-driven paradigms   (Tab. 3 P5-P8) for exploration goals. The should also consider the intent type as C3 mentions. Crucially, the twelve paradigms are not mutually exclusive: they can be flexibly combined and chained within a single application, enabling systems to shift fluidly between scaffolding, refinement, and repurposing workflows (Gmeiner et al., 2025; Wang et al., 2024d). Effective interfaces thus treat paradigms as composable building blocks rather than rigid templates, supporting diverse user needs. For example, IntentTagger (Gmeiner et al., 2025) introduces small, atomic intent tags enabling micro-prompting and region-level edits, illustrating chaining and covering multiple paradigms (e.g., P2, P4, P5, P7, P9). Reusing and chaining paradigms are not the end. Comparing the summarized paradigms and the entire IAI model, we can notice there are plenty of potential atomic paradigms that have not been explored. In the future, interaction designers may start from existing paradigms to design new ones for innovation, indicating the model’s generative power.
@@ 5.3. Usage Scenario 3: Applying the IAI model to Emerging Scenarios
Using the IAI model, the system designer begins by identifying the relevant entities: Human (H), Text Prompt (T), Generative AI (G), and Interaction (I), where Generative AI is to understand the user’s fine-tuning instruction and refine the initial prompt in the multi-agent system. Next, the system designer specifies the relations required to fulfill the user’s intent. First, the human issues the request (H $\rightarrow$ T). In this multi-agent system, producing a data video involves multiple tasks such as extracting insights, generating visualizations, crafting narration, and creating animations or annotations. Accordingly, a single instruction may map to different agent actions, such as the Designer agent animating Amazon’s line or the Data Analyst agent focusing on Amazon’s insights to craft narration. To ensure agent-specific accuracy, GenAI decomposes the text prompt and proposes candidate actions by different agents (T $\rightarrow$ G, G $\rightarrow$ I), and the human then reviews and confirms these suggestions (H $\rightarrow$ I). Thus, at this scenario, the interaction occurs after calling GenAI (Sec. 4.3-C1) and involves a text-only starting resource (Sec. 4.3-C2), consistent with the “Post-invocation, Prompt-only” category (Sec. 4.2.3). Finally, the confirmed interaction forms a new augmented instruction (I $\rightarrow$ Aug), which is passed back for another round of GenAI calling (chained by another follow-up paradigm).
@@ 5.4. Usage Scenario 4: Deriving New Atomic Paradigm Graphs
Consider the paradigm P11, Artifact-driven Prompt Enhancement (Fig. 6-P11), where GenAI initiates interactions based on information from artifacts (contextual information) and human prompts. Contextual information is critical in many scenarios, such as embodied AI in everyday life. But instead of the human initiating the conversation based on context like P11 (starting from H $\rightarrow$ T), what if the AI initiates the conversation proactively, as illustrated in Fig. 10 (left) and exemplified in Fig. 6-P9 and P10 (starting from A $\rightarrow$ G)?
@@ 6.1. Reflection
How should we apply interaction-augmented instructions? Adopting interaction-augmented instructions implies several foundational shifts in human–AI practice. First, expertise will shift from prompt engineering to instruction design. Users should compose reusable and multimodal instructions rather than optimizing text alone, while tools need to expose these capabilities with clear discoverability to support effective use (Srinivasan et al., 2019). Second, cognitive load and UX trade-offs must be managed: interactions can reduce text specification but add interface complexity, so designers should prioritize low-friction affordances (e.g., defaults, progressive disclosure, previews) to minimize overhead (Subramonyam et al., 2024). Third, evaluation can broaden beyond artifact quality to include axes that follow directly from the IAI distinction, such as correctness of local changes tied to interactions (referential fidelity), rounds to satisfactory output (convergence cycles), ease of steering behavior (controllability), and user comprehension of influences (provenance clarity).
@@
What can our model contribute to interaction-augmented instructions? The IAI model contributes to both engineering practice and HCI theory-building.   It provides greater descriptive, discriminative, and generative power than methods in prior work, such as Gao et al.  (Gao et al., 2024) and Shen et al.  (Shen et al., 2025a) by introducing two innovations. First, the model explicitly encodes information flow in the paradigm graph. Each edge denotes a precise link between entities, making differences in timing, source, and composition of instructions formally visible. Another key contribution is treating Aug as the explicit instruction to GenAI for the expressiveness of the model. Without it, designs that combine text prompts, interactions, and artifacts in different ways collapse into the broad “non-text interactions” category. It cannot systematically distinguish important differences, such as that between P1 and P2. With Aug, these differences become structurally visible, varying how entities feed into GenAI, and in what order, systematically yielding distinct paradigms. This mechanism directly enables both our comparative taxonomy and its generative extension, where new interfaces emerge from principled recombinations of information flow patterns.
@@
For practitioners, paradigm graphs serve as a design language that makes implicit design trade-offs explicit. Developers can compare alternatives, identify unexplored regions of the design space, and refine existing interfaces by recombining atomic paradigms. This resonates with prior calls to move beyond ad-hoc demonstrations toward systematic design frameworks in HCI (Beaudouin-Lafon, 2004). For researchers, our model complements prior taxonomies (Gao et al., 2024; Hu et al., 2025) and principle-based approaches (Riche et al., 2025) by introducing a formal representational structure that captures both entities and relations in human–GenAI workflows. For example, the three principles proposed in AI-Instruments (Riche et al., 2025) can be reinterpreted in IAI as properties of interaction flows rather than standalone design goals. Reification corresponds to whether and how intent is externalized into explicit intermediate representations (most notably Aug) that persist beyond a single interaction and can be inspected, reused, or chained. Reflection maps to interaction flows in which multiple alternatives are explicitly surfaced and negotiated over time, often through post-invocation paradigms where model outputs serve as anchors for iterative clarification. Grounding is realized through paradigms that bind instructions to concrete scopes (e.g., selected artifacts, regions, or prior outputs), thereby constraining interpretation through reference rather than description. Rather than treating these principles as orthogonal design heuristics, IAI shows how they manifest repeatedly across different atomic paradigms depending on interaction timing, available resources, and control flow. This shift reframes human–GenAI interaction from early-stage guideline analysis toward the concrete design and enactment of interaction structures, thereby enabling both actionable system designs and principled exploration of new paradigms. Such unified formalization enables cumulative comparison across systems and provides a substrate for analytical and generative methods, similar to how prior models advanced earlier eras of HCI (Card et al., 1986; Beaudouin-Lafon, 2000; Card et al., 1999).

### v3

@@ 4.3. Cross-paradigm Insights and Suggestions
In the previous section, we illustrate paradigm details and discuss their distinct design affordances based on recurring graph topologies derived from our IAI model. These paradigms and related findings demonstrate our model’s meaningful descriptive and discriminative powers to distinguish interface designs that prior frameworks conflate. For example, the motivating cases (Fig. 2 B–F) are clearly mapped to distinct paradigms (P1, P3, P4, P7, P9). We consider one critical resource of the powers is the IAI model’s information-flow structure. The structure allows us to use interaction timing and artifact availability as two principal axes to differentiate paradigms.
Starting from interaction timing (C1) and artifact availability (C2), this section provides concrete, model-grounded considerations that help designers decide which paradigm class aligns with design situations, as shown in Tab. 3 (C3), and how the IAI model and derived paradigms can be used to generate new interfaces (C4).
@@
C2. Use artifact availability (prompt-only vs. artifact-grounded) to select how intent should be expressed. Instructions in natural language alone can be ambiguous (e.g., make the flower brighter” without specifying which flower). The presence of an artifact shifts intent expression from purely linguistic descriptions to concrete, grounded references (Tab. 3-P4, P9-P12). By enabling users to select, highlight, or annotate artifact fragments, systems can reduce ambiguity and provide a stable foundation for commands (Singh et al., 2024; Chen et al., 2023). The consideration also aligns with design principles for AI-instruments by Riche et al.  (Riche et al., 2025), where they proposed that text prompts should be grounded in other artifacts. Designers should identify whether their system will receive an initial artifact; if so, artifact-grounded paradigms offer lower ambiguity, richer referential precision, and reduced prompting burden.
@@
C3. Within a chosen class, refine toward a specific paradigm based on how intent should be encoded. In these design paradigms, a key design choice is how user intent is materialized to augment the original text prompt. The first way is to directly edit or extend the text prompt with interactions (Tab. 2-P1, P5). An example is DreamSheets (Almeda et al., 2024), where users can combine prompt fragments easily with interactions. It provides a simple and direct way to improve original prompts but lacks the power to cater complex intent, such as non-linguistic or referential intent. To handle more complex intent, an approach is to leverage interactions to introduce additional non-linguistic information to text and form augmented instructions, such as structural or parametric information (Tab. 2-P2, P3, P6, P7, P9 and P11). For example, CoPrompt (Feng et al., 2024) allows users to drag and drop prompt fragments to form a multi-level list of instructions. It materializes the non-linguistic intent expression in a convenient way with suitable interaction design. Regarding the referential intent, the key interaction is to link text prompts with artifacts and generate augmented instructions (Tab. 2-P4, P10). It addresses the challenges in describing the linkage between texts and artifacts by materializing users’ intent with direct interactions on artifacts. Lastly, a unique case is to generate new interaction widget by text prompts and facilitate intent expression through direct manipulation (Tab. 2-P8, P12). GenAI creates reusable and persistent interactive widgets to materialize the meta intent and allows follow-up similar intent expression with simple interactions with widgets. A notable example is DynaVis (Vaithilingam et al., 2024). The meta intent like changing visual element colors can be materialized as a widget. The specific color to use can be directly selected by interactions. The summarized design cheat sheet mapping common design situations to recommended paradigms is shown in Tab. 3. These differences correspond to distinct subgraph configurations in the atomic paradigm graphs. Designers can therefore match their intent-expression needs (e.g., decomposition, parameterization, selection specificity) to the paradigm whose graph topology supports that form of control.
@@
Table 4. Operationalizing Generative Power: Potential methods for designers and researchers to apply the IAI model and derived paradigms. In this table, entry points refer to whether the designer or researcher should look into IAI model or our summarized paradigms at the beginning of the usage method.
@@
C4. Reusing, chaining, and innovating atomic paradigms for adapting interaction design to new scenarios. Our summarized paradigms and considerations C1-C3 above provide a concise mapping from task requirements to interface affordances. Designers can start from intent clarity and artifact availability to select appropriate paradigms for interface implementation. For example, they can use the pre-invocation, artifact-grounded paradigm (Tab. 3-P5) for tasks that start from existing artifacts and demand precision. They can apply post-invocation, prompt-driven paradigms (Tab. 3 P5-P8) for exploration goals. The should also consider the intent type as C3 mentions. Crucially, the twelve paradigms can be flexibly combined and chained within a single application, enabling systems to shift fluidly between scaffolding, refinement, and repurposing workflows (Gmeiner et al., 2025; Wang et al., 2024d). Effective interfaces thus treat paradigms as composable building blocks rather than rigid templates, supporting diverse user needs. For example, IntentTagger (Gmeiner et al., 2025) introduces small, atomic intent tags enabling micro-prompting and region-level edits, illustrating chaining and covering multiple paradigms (e.g., P2, P4, P5, P7, P9). Reusing and chaining paradigms are not the end. Comparing the summarized paradigms and the entire IAI model, we can notice there are plenty of potential atomic paradigms that have not been explored. In the future, interaction designers may start from existing paradigms to design new ones for innovation, indicating the model’s generative power.
@@ 5.3. Usage Scenario 3: Applying the IAI model to Emerging Scenarios
Using the IAI model, the system designer begins by identifying the relevant entities: Human (H), Text Prompt (T), Generative AI (G), and Interaction (I), where Generative AI is to understand the user’s fine-tuning instruction and refine the initial prompt in the multi-agent system. Next, the system designer specifies the relations required to fulfill the user’s intent. First, the human issues the request (H $\rightarrow$ T). In this multi-agent system, producing a data video involves multiple tasks such as extracting insights, generating visualizations, crafting narration, and creating animations or annotations. Accordingly, a single instruction may map to different agent actions, such as the Designer agent animating Amazon’s line or the Data Analyst agent focusing on Amazon’s insights to craft narration. To ensure agent-specific accuracy, GenAI decomposes the text prompt and proposes candidate actions by different agents (T $\rightarrow$ G, G $\rightarrow$ I), and the human then reviews and confirms these suggestions (H $\rightarrow$ I). Thus, at this scenario, the interaction occurs after calling GenAI (Sec. 4.3-C1) and involves a text-only starting resource (Sec. 4.3-C2), consistent with the “Post-invocation, Prompt-only” category (Sec. 4.2.3). Finally, the confirmed interaction forms a new augmented instruction (I $\rightarrow$ Aug), which is passed back for another round of GenAI calling (chained by a follow-up paradigm).
@@ 5.4. Usage Scenario 4: Deriving New Atomic Paradigm Graphs
Consider the paradigm P11, Artifact-driven Prompt Enhancement (Fig. 6-P11), where GenAI initiates interactions based on information from artifacts (contextual information) and human prompts. Contextual information is critical in many scenarios, such as embodied AI in everyday life. But instead of a human initiating the conversation based on context like P11 (starting from H $\rightarrow$ T), what if the AI initiates the conversation proactively, as illustrated in Fig. 10 (left) and exemplified in Fig. 6-P9 and P10 (starting from A $\rightarrow$ G)?
@@ 6.1. Reflection
How should we apply interaction-augmented instructions? Adopting interaction-augmented instructions implies several foundational shifts in human–AI practices. First, expertise will shift from prompt engineering to instruction design. Users should compose reusable and multimodal instructions rather than optimizing text alone, while tools need to expose these capabilities with clear discoverability to support effective use (Srinivasan et al., 2019). Second, cognitive load and UX trade-offs must be managed: interactions can reduce text specification but add interface complexity, so designers should prioritize low-friction affordances (e.g., defaults, progressive disclosure, previews) to minimize overhead (Subramonyam et al., 2024). Third, tool evaluation can broaden beyond artifact quality to include axes that follow directly from the IAI distinction, such as correctness of local changes tied to interactions (referential fidelity), rounds to satisfactory output (convergence cycles), ease of steering behavior (controllability), and user comprehension of influences (provenance clarity).
@@
What can our model contribute to interaction-augmented instructions? The IAI model contributes to both engineering practice and HCI theory-building. It provides greater descriptive, discriminative, and generative power than methods in prior work, such as Gao et al.  (Gao et al., 2024) and Shen et al.  (Shen et al., 2025a), by introducing two innovations. First, the model explicitly encodes information flow in the paradigm graph. Each edge denotes a precise link between entities, making differences in timing, source, and composition of instructions formally visible. Another key contribution is treating Aug as an explicit instruction to GenAI for the expressiveness of the model. Without it, designs that combine text prompts, interactions, and artifacts in different ways collapse into the broad “non-text interactions” category. It cannot systematically distinguish important differences, such as that between P1 and P2. With Aug, these differences become structurally visible, varying how entities feed into GenAI, and in what order, systematically yielding distinct paradigms. This mechanism directly enables both our comparative taxonomy and its generative extension, where new interfaces emerge from principled recombinations of information flow patterns.
@@
For practitioners, paradigm graphs serve as a design language that makes implicit design trade-offs explicit. Developers can compare alternatives, identify unexplored regions of the design space, and refine existing interfaces by recombining atomic paradigms. This resonates with prior calls to move beyond ad-hoc demonstrations toward systematic design frameworks in HCI (Beaudouin-Lafon, 2004). For researchers, our model complements prior taxonomies (Gao et al., 2024; Hu et al., 2025) and principle-based approaches (Riche et al., 2025) by introducing a formal representational structure that captures both entities and relations in human–GenAI workflows. For example, the three principles proposed in AI-Instruments (Riche et al., 2025) can be reinterpreted in IAI as properties of interaction flows rather than standalone design goals. Reification corresponds to whether and how intent is externalized into explicit intermediate representations (most notably Aug) that persist beyond a single interaction and can be inspected, reused, or chained. Reflection maps to interaction flows in which multiple alternatives are explicitly surfaced and negotiated over time, often through post-invocation paradigms where model outputs serve as anchors for iterative clarification. Grounding is realized through paradigms that bind instructions to concrete artifact scopes (e.g., selected artifacts, regions, or prior outputs), thereby constraining interpretation through reference rather than description. Rather than treating these principles as orthogonal design heuristics, IAI shows how they manifest repeatedly across different atomic paradigms depending on interaction timing, available resources, and control flow. This shift reframes human–GenAI interaction from early-stage guideline analysis toward the concrete design and enactment of interaction structures, thereby enabling both actionable system designs and principled exploration of new paradigms. Such unified formalization enables cumulative comparison across systems and provides a substrate for analytical and generative methods, similar to how prior models advanced earlier eras of HCI (Card et al., 1986; Beaudouin-Lafon, 2000; Card et al., 1999).